OpenStudy (anonymous):

Can someone let me know if my answer is correct? I can provide the answers if needed. 1. Describe how the graph of y=|x| and y=|x| - 15 are related? My answer: The two graphs have the same y-intercept. The second graph is steeper than y=|x|.

OpenStudy (anonymous):

If the graphs have the same y intercept, this means that when x=0 the y values are the same. Are they the same in this case?

OpenStudy (anonymous):

So the answer would be the two graphs are the same in this case, correct? @tom982

OpenStudy (anonymous):

No, just think about it :) Regardless of the value of x we put into the equation, the y value is always going to be 15 units lower in y=|x|-15 than it is in y=|x|.

OpenStudy (solomonzelman):

The absolute value would follow the same rule for what is inside and outside the parenthesis. \(\large\color{ blue }{\large {\bbox[5pt, lightyellow ,border:2px solid white ]{ \large\text{ }\\ \begin{array}{|c|c|c|c|} \hline~~~~~~~~~~~~~~~~~~~~~~~~~~~\textbf{Shifts}~~~~~~~~~~~~~~~~~~~~~~~~~~~&~\bf{c~~~units~~~~} \\ \hline \\f(x)= ∜x ~~~ ⇒ ~~~ f(x)= \sqrt[4]{x \normalsize\color{red }{ -~\rm{c}} } &~\rm{to~~the~~right~} \\ \text{ } \\ f(x)= ∜x ~~~ ⇒ ~~~ f(x)= \sqrt[4]{x \normalsize\color{red}{ +~\rm{c}} } &~\rm{to~~the~~left ~} \\ \text{ } \\ f(x)= ∜x ~~~ ⇒ ~~~ f(x)= ∜x \normalsize\color{red}{ +~\rm{c} } &~\rm{up~} \\ \text{ } \\ f(x)= ∜x ~~~ ⇒ ~~~ f(x)= ∜x \normalsize\color{red}{ -~\rm{c} } &~\rm{down~} \\ \\ \hline \end{array} }}}\)

OpenStudy (solomonzelman):

dang question marks. Hell no!

OpenStudy (anonymous):

Thanks @tom982 I believe I have the answer, sorry for my slowness :) Thank you lol @SolomonZelman

OpenStudy (solomonzelman):

Anytime... I am also slow... everyone one is slow in one subject or the other:)

OpenStudy (anonymous):

@Brielani, no problem, we were all beginners once. Glad you understand it though.